On January 1, 2020, NoDice Corporation issues $540,000, 5-year, 12% bonds for $529,000. Interest is paid semiannually on January 1 and July 1. NoDice Corporation uses the straight-line method of amortization. The company's fiscal year ends on December 31. The amount of bond interest expense on July 1, 2020 is: 

$31,300 

$33,500 

$32,400 

$65,900 
Answer #1

Discount on issue of bonds = $540000 - 529000 = $11000

Interest paid in cash = $540000 x 12% x 6/12 = $32400

Bond Interest Expense = $32400 + 11000/10 = $33500

Answer is b. $33500
Answer #2

ANSWER :


Discount on bond = 540000 - 529000 = 11000 ($)


Proportion of discount distributed over 6 month period (unto July 20) :

= 11000 / (5 years maturity * 2 periods of 6 months per year)

= 1100 ($)


Interest for 6 months on bond to be paid on July 20 

= 540000 * 12/100 * (6 month / 12 months per year)

= 32400 ($)


So, effective interest on. Bond 

= Interest on the bond + adjustment of discount

= 32400 + 1100

= 33500 ($) : 2nd Option (ANSWER).

answered by: Tulsiram Garg